Samantha deposits $2,000 in a saving account that pays an annual interest rate of 3 percent. Though the expected inflation was 2%, over the course of the year the inflation rate is 4 percent. Calculate how much money Samantha has at the end of the year and how much her purchasing power has changed.
1
Expert's answer
2020-11-16T07:16:52-0500

Amount at the end of the year = 2000 + 3"\/" 100 "\times" 2000

= 2000 + 60

= "\$" 2060

Purchasing power when inflation rate was 2 "\%"

= 2060-2"\%"

= "\$" 2018.8

Purchasing power when inflation rate actually was 4 "\%"

= 2060-4"\%"

="\$" 1977.6

Hence as the result of the rise in the actual rate of inflation , the purchasing power dropped by "\$" 41.2